OpenStudy (anonymous):

If a=3^p, b=3^q, c=3^r, and d=3^s and if p,q,r,and s are positive integers, determine the smallest value of p+q+r+s such that: a^2+b^3+c^5=d^7

hartnn (hartnn):

interesting.

hartnn (hartnn):

a=3^p --> p = log[3] a p+q+r+s = log[3] abcd

OpenStudy (anonymous):

log?

OpenStudy (anonymous):

i kind of understand the first part but i am not quite sure about the second part p+q+r+s = log[3] abcd

hartnn (hartnn):

a^2+b^3 +c^5 = d^7 gives 3^{2p} + 3^ {3q} + 3^ {5r} = 3^ {7s} oh, i have used that log x + log y = log xy

hartnn (hartnn):

i am just trying, idk the solution

OpenStudy (anonymous):

it's ok, we should try any attempt

hartnn (hartnn):

but now did u get this p+q+r+s = log[3] (abcd) ?

OpenStudy (anonymous):

yep

hartnn (hartnn):

is this calculus related ? which topic does this question belong ?

OpenStudy (anonymous):

no, it is not caculus related. this is an exponent question

OpenStudy (anonymous):

Well d must be odd which means s must also be odd.

OpenStudy (anonymous):

why does d must be odd?

OpenStudy (turingtest):

*

OpenStudy (anonymous):

d must be odd because